The magnetic field at the centre of a solenoid
of length 2.5 m and diameter 0.8 m is
0.5 tesla. Calculate the energy stored in the
field of the solenoid.
1
Expert's answer
2019-05-06T10:42:51-0400

The energy

"W=\frac{B^{2}L\pi D^{2}}{8\mu_0}=\frac{0.25\cdot 2.5\cdot 0.64 }{32\cdot 10^{-7}} = 130 (kJ)"